OpenStudy (anonymous):

Which of the equations describes a line on a graph with the points (0,2) and (1,1)? A. y = x – 2 B. y = x + 2 C. y = -x + 2 D. y = –x – 2

OpenStudy (lgbasallote):

Just use two point form y - 2 = [(1-2)/(1)](x) y - 2 = -x y = -x +2 ANswer: C